Calculate the price of a zero-coupon bond that matures in 12 years if the market interest rate is 6.05 percent. Assume semiannual compounding.
Answer #1

F Zero Coupon Bond Value (1r) F = face value of bond т 3 rate or yield t time to maturity

n = 12 * 2 = 24 semi-annual periods, r = 6.05%/2 = 3.025% (semi-annually)

1000 Value (10.03025)24

Value = 1000/2.0447

Value = $489.08